# Disentangling the Impact of Quasiparticles and Two-Level Systems on the Statistics of Superconducting Qubit Lifetime

### v1

@@ Abstract
with $\eta(T)=\frac{16E_{\text{J}}}{\hbar\pi}\sqrt{\frac{E_{\text{C}}}{8E_{\text{J}}}}\sqrt{\frac{2\Delta}{\pi k_{\text{B}}T}}\exp\left(\frac{\hbar\omega_{\text{q}}}{2k_{\text{B}}T}\right)K_{0}\left(\frac{\hbar\omega_{\text{q}}}{2k_{\text{B}}T}\right)$. Here, the non-equilibrium and equilibrium QPs are assumed to be independent [28]. In Eq. (2), we define $V^{\text{0}}_{\text{eff}}$ and $V^{\text{th}}_{\text{eff}}$ as the effective volumes for non-equilibrium and thermal QPs, respectively, to reflect their distinct contributions to the fluctuations. Using previously obtained $x^{0}_{\text{qp}}$, $x^{\text{th}}_{\text{qp}}$, and $\Delta$ [see Table 1], we can fit the experimentally extracted $\sigma^{2}_{\text{qp}}(\Gamma_{1})$ and obtain the values of $V^{\text{0}}_{\text{eff}}$ and $V^{\text{th}}_{\text{eff}}$ for each qubit. The fitted curves are shown as solid lines in Fig. 3(a), and the fitted $V^{\text{0}}_{\text{eff}}$ and $V^{\text{th}}_{\text{eff}}$ are also listed in Table 1. In the non-equilibrium regime, the effective volume is expected to be proportional to $(v_{\text{F}}\cdot\tau_{\text{r}})^{3}$, with the Fermi velocity $v_{\text{F}}$ being the same for all qubits, and $\tau_{\text{r}}$ the QP recombination rate that is inversely proportional to $x_{\text{qp}}^{0}$ [18]. Therefore, we have $V^{0}_{\text{eff}}\propto(1/x_{\text{qp}}^{0})^{3}$. We notice that the ratio of $V^{0}_{\text{eff}}$ between device C (B) and A is about 13 (5) from the fitting. Indeed, this aligns with our data: the ratio of $x_{\text{qp}}^{0}$ between device A and C (B) is about 2.5 ($2.5^{3}\sim 15$). Moreover, the difference in $x_{\text{qp}}^{0}$ can be supported by the diffusion process, from which the $x_{\text{qp}}^{0}$ ratio is calculated to be 2.7 between different pad sizes. In the thermal regime, the QP density dominated by the thermal generation in the Al junction leads increases significantly but does not substantially suppress the $V_{\text{eff}}^{\text{th}}$. On the other hand, $V_{\text{eff}}^{\text{th}}$ cannot exceed the upper limit set by the non-equilibrium QPs from the qubit pads when $k_{\text{B}}T$ approaches the Al gap energy. Therefore we approximate the $V_{\text{eff}}^{\text{th}}$ as a constant in Eq. (2). From the fit, we find the effective volume $V^{\text{th}}_{\text{eff}}$ for all the qubits approaches the same value, close to the volume of the Josephson junction $\sim 0.013\,\mu\textrm{m}^{3}$. This is as expected since all qubits consist of the same Al junctions. Here we define the junction volume as the product of the junction area and the total thickness of Al layers.
@@
Qubit $x_{\text{qp}}^{0}$ $\Delta/2\pi$ (GHz) $V_{\text{eff}}^{0}$ ($\mu\text{m}^{3}$) $V_{\text{eff}}^{\text{th}}$ ($\mu\text{m}^{3}$) $\Gamma_{\text{tls}}$ (MHz)
@@
In Fig. 3(b), we plot the $\sigma^{2}_{\text{TLS}}(\Gamma_{1})$ induced by TLSs. Compared to the variance from QPs, the $\sigma^{2}_{\text{TLS}}(\Gamma_{1})$ here show an overall weak dependence on temperature, with some local features. It is also observed that $\sigma^{2}_{\text{TLS}}(\Gamma_{1})$ of qubit A is larger than qubit B and C, while the latter two show some consistency across the temperature range. We suggest that this can be qualitatively explained by TLS theory as well. The variance of qubit depolarization due to the interaction between a single TLS and an assemble of TLFs is given by [27]:
@@
where $A$ denotes a constant related to the coupling strength and qubit matrix element, $\gamma_{2}$ describes the linewidth of the TLS, and $g_{i}$ the coupling strength between the TLS and the $i$th TLF, with frequency $\omega_{\text{t,i}}$. When the temperature increases, the TLS linewidth gets larger, suppressing the amplitude of the Lorentzian, and thus the variance of the fluctuation. On the other hand, higher temperature leads to more frequent switching of the TLFs, increasing the last factor in Eq. (3). This discussion is based on a single TLS. In practice, multiple TLSs can couple simultaneously to the qubit, each with different dephasing rates and frequencies. As a result, the above two competing factors for different TLSs may lead to the observed behavior in Fig. 3(b).
Our work underscores a practical need to stabilize qubit lifetimes. On one hand, theoretical proposals based on driving the qubits [36] or the TLSs [27] could provide ways to stabilize the fluctuations due to TLSs, without affecting the average lifetime. On the other hand, protocols for reducing fluctuations from QPs remain to be developed but may benefit from strategies to reduce QPs in general [34, 37, 38, 32].
In summary, we perform prolonged, statistical $T_{1}$ measurements on three superconducting transmon qubits across a broad temperature range. Different geometrical footprints and superconducting materials of these qubits enable us to disentangle the impacts of QPs and TLSs on the average and fluctuation of the qubit lifetime. The improvements in the average qubit lifetime can be explained by an optimization of the qubit design as well as a reduced loss of surface dielectric. More importantly, we find that the QP densities near the Josephson junction depend on the qubit footprints, i.e., higher density in the small-footprint qubit. We further propose the concept of effective QP volume that scales cubically with the inverse of the non-equlibrium QP density, therefore providing quantitative support in understanding the discrepancy of the lifetime fluctuations in the qubits with different footprints. We suggest that these findings can provide a guideline for future optimization of qubit design and engineering.

### v2

@@ Abstract
with $\eta(T)=\frac{16E_{\text{J}}}{\hbar\pi}\sqrt{\frac{E_{\text{C}}}{8E_{\text{J}}}}\sqrt{\frac{2\Delta}{\pi k_{\text{B}}T}}\exp\left(\frac{\hbar\omega_{\text{q}}}{2k_{\text{B}}T}\right)K_{0}\left(\frac{\hbar\omega_{\text{q}}}{2k_{\text{B}}T}\right)$. Here, the non-equilibrium and equilibrium QPs are assumed to be independent [30]. In Eq. (2), we define $V^{\text{0}}_{\text{eff}}$ and $V^{\text{th}}_{\text{eff}}$ as the effective volumes for non-equilibrium and thermal QPs, respectively, to reflect their distinct contributions to the fluctuations. Using previously obtained $x^{0}_{\text{QP}}$, $x^{\text{th}}_{\text{QP}}$, and $\Delta$ [see Table 1], we can fit the experimentally extracted $\sigma^{2}_{\text{QP}}(\Gamma_{1})$ and obtain the values of $V^{\text{0}}_{\text{eff}}$ and $V^{\text{th}}_{\text{eff}}$ for each qubit. The fitted curves are shown as solid lines in Fig. 3(a), and the fitted $V^{\text{0}}_{\text{eff}}$ and $V^{\text{th}}_{\text{eff}}$ are also listed in Table 1. In the non-equilibrium regime, the effective volume is expected to be proportional to $(v_{\text{F}}\cdot\tau_{\text{r}})^{3}$, with the Fermi velocity $v_{\text{F}}$ being the same for all qubits, and $\tau_{\text{r}}$ the QP recombination rate that is inversely proportional to $x_{\text{QP}}^{0}$ [19]. Therefore, we have $V^{0}_{\text{eff}}\propto(1/x_{\text{QP}}^{0})^{3}$. We notice that the ratio of $V^{0}_{\text{eff}}$ between device C (B) and A is about 13 (5) from the fitting. Indeed, this aligns with our data: the ratio of $x_{\text{QP}}^{0}$ between device A and C (B) is about 2.5 ($2.5^{3}\sim 15$). Moreover, the difference in $x_{\text{QP}}^{0}$ can be supported by the diffusion process, from which the $x_{\text{QP}}^{0}$ ratio is calculated to be 2.7 between different pad sizes. In the thermal regime, the QP density dominated by the thermal generation in the Al junction leads increases significantly but does not substantially suppress the $V_{\text{eff}}^{\text{th}}$. On the other hand, $V_{\text{eff}}^{\text{th}}$ cannot exceed the upper limit set by the non-equilibrium QPs from the qubit pads when $k_{\text{B}}T$ approaches the Al gap energy. Therefore we approximate the $V_{\text{eff}}^{\text{th}}$ as a constant in Eq. (2). From the fit, we find the effective volume $V^{\text{th}}_{\text{eff}}$ for all the qubits approaches the same value, close to the volume of the Josephson junction $\sim 0.013\,\mu\textrm{m}^{3}$, which is as expected since all qubits consist of the same Al junctions. Here, we define the junction volume as the product of the junction area and the total thickness of Al layers.
@@
Qubit $x_{\text{QP}}^{0}$ $\Delta/2\pi$ (GHz) $V_{\text{eff}}^{0}$ ($\mu\text{m}^{3}$) $V_{\text{eff}}^{\text{th}}$ ($\mu\text{m}^{3}$) $\Gamma_{\text{TLS}}$ ($\mu\text{s}^{-1}$)
@@
In Fig. 3(b), we plot the $\sigma^{2}_{\text{TLS}}(\Gamma_{1})$ induced by TLSs. Compared to the variance from QPs, the $\sigma^{2}_{\text{TLS}}(\Gamma_{1})$ here show an overall weak dependence on temperature, with some local features. It is also observed that $\sigma^{2}_{\text{TLS}}(\Gamma_{1})$ of qubit A is notably larger than qubits B and C, while the latter two show some consistency in the temperature range. We suggest that this can be qualitatively explained by TLS theory as well. The variance of qubit depolarization due to the interaction between a single TLS and an assemble of TLFs is given by [29]:
@@
where $A$ denotes a constant related to the coupling strength and qubit matrix element, $\gamma_{2}$ describes the linewidth of the TLS, and $g_{i}$ the coupling strength between the TLS and the $i$th TLF, with frequency $\omega_{\text{t,i}}$. When the temperature increases, the TLS linewidth gets larger, suppressing the amplitude of the Lorentzian and, thus, the variance of the fluctuation. On the other hand, higher temperature leads to more frequent switching of the TLFs, increasing the last factor in Eq. (3). This discussion is based on a single TLS. In practice, multiple TLSs can couple simultaneously to the qubit, each with different dephasing rates and frequencies. As a result, the above two competing factors for different TLSs may lead to the observed behavior in Fig. 3(b) [23].
Our work underscores a practical need to stabilize qubit lifetimes. On one hand, theoretical proposals based on driving the qubits [38] or the TLSs [29] could provide ways to stabilize the fluctuations due to TLSs without affecting the average lifetime. However, protocols to reduce QP fluctuations have yet to be developed, but may benefit from strategies to reduce QPs in general [36, 39, 40, 34].
We perform prolonged, statistical $T_{1}$ measurements on three superconducting transmon qubits across a broad temperature range. Different geometrical footprints and superconducting materials of these qubits enable us to disentangle the impacts of QPs and TLSs on the average and fluctuation of the qubit lifetime. The improvements in the average qubit lifetime can be explained by an optimization of the qubit design as well as a reduced loss of surface dielectric. More importantly, the QP densities near the Josephson junction depend on the qubit footprints, i.e., higher density in the small-footprint qubit. We further propose the concept of effective QP volume, which scales cubically with the inverse of the non-equilibrium QP density and provides quantitative support for understanding the discrepancy between the lifetime fluctuations in qubits with different footprints. We suggest that these findings can guide future qubit design and engineering optimization.